Rafael owned an apartment building that burned down. The empty lot is worth $70,000 and Rafael has received $200,000 from the insurance company. Rafael plans to build another apartment building that will cost $275,000. His real estate adviser estimates that the expected value of the finished building on the real estate market will be $385,000 next year. The discount/interest rate is 10%? What are the NPV and IRR of this decision?
1
Expert's answer
2013-07-16T12:02:15-0400
PV of a finished building PV=385000/1.1=350000. So, NPV of the project NPV=350000-275000=75000.
To calculate IRR we solve an equation 275000=385000/(1+r),r=(85000-275000)/275000=40%.